from a group of 6 men and 8 women a committee of 3 is to be selected at random find the probability that all 3 are men or all 3 woman
Answer by chana(69)

from a group of 6 men and 8 women a committee of 3 is to be selected at random find the probability that all 3 are men or all 3 woman..
sol :
p(e) = (n(e))/(n(s))
n(s) = 3 mem from 14 people
14c3 = 364
n(e) = 3 from total men and 3 from total women
= 6c3 + 8c3
= 76
so p(e) = 76/364
=19/91.